# AI-Slop to AI-Polish? Aligning Language Models through Edit-Based Writing Rewards and Test-time Computation

### v2

@@ Appendix A Appendix
Appendix B Example LAMP Sample
@@
Appendix C Generative WRQM Prompt Formats
@@
Appendix D Example Rationales For Generative WQRM Training
@@ D.1 Datasets
Appendix E Example Human Mimic Samples
@@
Appendix F Example COT Editing Prompt
@@
Appendix G Expert Annotation Result Breakdown
@@
Appendix H Comparison
Table 9 shows 3 different versions of the same paragraph. First Draft along with edited versions (Random and Best Edit) with respect rewards from WQRM. Experts rank this triplet as Best Edit $>$ Random Edit $>$ First Draft.
@@ Appendix H Comparison
Appendix I Expert Annotation Interface
@@
World with his new family. The narrator declined, fabricating an excuse about school, despite

### v3

@@ Appendix A Appendix
Which is better quality writing?
A.1 Example LAMP Sample
@@
A.2 Generative WRQM Prompt Formats
@@
A.3 Example Rationales For Generative WQRM Training
@@ A.4 Datasets
A.5 Writing Quality Benchmark Difficulty Analysis
Figure 5: Gap Analysis of WQ datasets leveraging the WQRM-PR model.
In order to understand the relative difficulty of the datasets within the WQ benchmark, we performed an analysis leveraging our trained WQRM model. For each sample (consisting of two writing samples with a known human preference), we computed the WQRM score for each sample, and compiled the result for each of the five datasets in WQRM. Figure 5 plots the average of the preferred vs. less-preferred scores on each dataset.
This analysis allows to make several observations. First, the average WQRM gap is directly proportional with model performance on the benchmark. The Synthetic Mirror dataset has the largest average gap according to WQRM-PR (2.4 on average), and we find that many models achieve very close to perfect performance (98%+) on this dataset. On the other hand, the gap (according to WQRM-PR) is very small on Style Mimic (0.12) and LMArena (0.02), which aligns with many models performing at or very slightly above chance on these datasets. Second, the absolute scores for the low and high samples are indicative of the origin of the samples. Style Mimic is the only dataset to include Human-Human comparisons (both written by professionals), and the scores of both the worse and better writing samples are high (7.57 and 7.69). LMArena has a similarly small gap, but achieved with lower pair scores (5.99 and 6.02). Third, we find that the WQ dataset includes a mix of high-gap (easy) and low-gap datasets. For low-gap samples, those can be with both having lower scores (two AI-generated samples), or two high-scoring samples (two human-written samples). This confirms the breadth of evaluation included in the WQ benchmark, which is a primary objective of the WQ benchmark.
We note that this analysis should be taken with a grain of salt: the WQRM-PR model is not a perfect score predictor, and is only a proxy for analysis, since true scores would require large-scale professional annotation (which is cost-prohibitive). But this analysis matches some expectations, and provides additional evidence of the proper calibration of the WQRM-PR model, and of the breadth of evaluation in the WQ benchmark.
A.6 Example Human Mimic Samples
@@
A.7 Example COT Editing Prompt
@@
A.8 Expert Annotation Result Breakdown
@@
A.9 Comparison
Table 9 shows 3 different versions of the same paragraph. First Draft along with edited versions (Random and Best Edit) with respect rewards from WQRM. Experts rank this triplet as Best Edit >> Random Edit >> First Draft.
@@ A.9 Comparison
A.10 Expert Annotation Interface
@@
World with hiss new family. The narrator declined, fabricating an excuse about school, despite
@@
A.11 Better Calibrated WQRM model for Content and Quality Experiment
Since WQRM was only trained on samples from LAMP, which consists of AI-generated paragraphs edited by MFA students, it doesn’t fully know how to reward higher-quality human writing. For this purpose, we added 100 paragraphs written by 5 award-winning authors (20 each) to our training data. We chose 5 authors who were part of the Style Mimic data. Each paragraph written by an award-winning author was assigned a score of 10.0. Even within writing from trained professionals, there is significant variability. To address this we source an additional 80 independent paragraphs written by MFA students published in prestigious literary magazines such as Electric Lit, Joyland, Paris Review and add to our training data. Each paragraph written by an MFA student was assigned a score of 7.5  [^7] This was a design decision where 5 is average and 10 is the best, and 7.5 is a mid-point.. Publication at a venue already means these paragraphs have undergone scrutiny and are of decent quality. After adding these 180 samples to LAMP-PR training set, we retrained WQRM.